# Elastic, Quasielastic, and Superelastic Electron Scattering from Thermal Lattice Distortions in Perfect Crystals

### v1

@@ VII.5 Planckian diffusion and linear-in-$T$ resistivity
arising from elastic momentum scrambling in a time-dependent lattice background. This form of Planckian diffusion does not rely on quasiparticles, strong inelastic scattering, or rapid energy dissipation; it follows directly from quantum dynamics in a fluctuating potential and has been demonstrated explicitly in wave-on-wave simulations of the full electron-lattice problem [5, 38].
@@
so that recasting $D$ in terms of an effective relaxation time yields
@@
up to factors of order unity. From this perspective, the Planckian timescale is not a fundamental microscopic bound, but an emergent parametrization of quantum-limited diffusion combined with equilibrium thermodynamics. It constrains how rapidly charge spreads in space, not how rapidly energy must be irreversibly transferred to internal degrees of freedom.
This viewpoint also clarifies the emergence of linear-in-$T$ resistivity across and below the Debye scale. Within a conventional Bloch-Grüneisen picture, the Debye temperature sets the scale at which all phonon modes become available for inelastic scattering. In contrast, within the quantum acoustics framework there is no special anomaly associated with crossing the Debye temperature. Linear-in-$T$ resistivity arises smoothly as coherence is lost in a time-dependent elastic background, while energy relaxation remains comparatively slow. The Debye scale therefore marks neither the onset of momentum relaxation nor a transition to a new inelastic regime, but rather a coherence crossover.
VII.6 Mott-Ioffe-Regel crossover as loss of quasiparticle coherence
The Mott-Ioffe-Regel (MIR) condition $k_{F}\ell\sim 1$ is often interpreted as a boundary beyond which quasiparticles “cease to exist” and transport must become anomalous. From the present perspective, this interpretation is misleading. The MIR condition marks the point at which a semiclassical mean free path loses a literal geometric meaning, not a dynamical singularity in transport.
Within the quantum acoustics framework, transport evolves smoothly through $k_{F}\ell\sim 1$. Diffusion persists, but it is no longer usefully described in terms of well-defined trajectories between discrete scattering events. Crucially, this breakdown of a quasiparticle picture does not coincide with a loss of coherence. Were strongly inelastic processes truly dominant at the MIR scale, interference phenomena such as weak localization and quantum oscillations would be rapidly suppressed, contrary to experimental observation. Instead, experiments show that momentum relaxation can be strong while inelastic energy relaxation remains weak.
Wave-on-wave simulations sharpen this picture by explicitly demonstrating that momentum can be transferred elastically to a thermally distorted lattice background while preserving electronic energy. The MIR crossover therefore reflects a loss of quasiparticle coherence rather than the onset of strong dissipation, fully consistent with elastic-dominated momentum relaxation in a dynamical lattice environment.
@@ VIII Summary
Summary
In the conventional clamped-lattice formulation of electron-phonon scattering, the factor $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$ associated with rigid translation of the lattice background is removed by assumption. This elimination suppresses an otherwise allowed elastic momentum-transfer channel, forcing momentum relaxation to be identified exclusively with inelastic phonon creation or annihilation. By restoring the lattice translational zero mode, we show that this identification is not required by symmetry, conservation laws, or kinematics. In a finite crystal, elastic pseudomomentum transfer between an electron and the lattice background is generically present and can dominate momentum relaxation at low and intermediate temperatures.
Making this elastic channel explicit requires abandoning several silent but consequential restrictions embedded in the traditional formulation. First, the lattice must be treated as a finite, isolated object so that its center-of-mass degree of freedom remains dynamical. (If the crystal is attached to a laboratory or substrate, this simply enlarges the background without altering the underlying momentum bookkeeping.) Second, discrete translational symmetry must be implemented symmetrically, with foreground and background pseudomomenta treated as equally real and conserved quantities. Third, the electron-lattice interaction must be retained in its full density-density form, rather than truncated in a way that projects out elastic momentum-transfer processes by construction. When these elements are retained, elastic scattering of electrons from thermal lattice distortions is not merely allowed but unavoidable.
In this framework, pseudomomentum is exchanged between the electron and the lattice background, while total mechanical momentum is conserved through recoil of the lattice zero mode, without the necessity of phonon creation. The long-standing identification of lattice motion with inelasticity is thus revealed to be an artifact of an incomplete theoretical description.
Restoring elastic momentum-transfer processes to their proper role provides a more complete and internally consistent foundation for electronic transport in crystals. It offers a natural interpretation of experimental observations that indicate robust elastic scattering in clean materials and clarifies how momentum relaxation can be efficient even when energy relaxation is slow. More broadly, these results raise the possibility that in at least some strange-metal regimes, momentum relaxation is dominated by elastic scattering in a time-dependent background, with genuinely inelastic processes playing a secondary role in establishing energy equilibration rather than setting the transport rate.
IX Acknowledgements
@@
Appendix A Elastic and inelastic scattering in a tight-binding, isolated one-dimensional chain
We consider a finite one-dimensional tight-binding chain of $N$ identical atoms, isolated in space, with nearest-neighbor springs and an electron moving in a basis of localized $p$ orbitals. The chain possesses an exact translational zero mode, corresponding to its center-of-mass (CM) motion. This allows the electron to exchange momentum with the lattice as a whole without necessarily exciting internal phonon modes. As a result, at finite temperature there exists a strictly elastic backscattering channel for an electronic wavepacket.
@@ Appendix A Elastic and inelastic scattering in a tight-binding, isolated one-dimensional chain
Let the atomic coordinates be written as
@@
where $R$ is the center-of-mass coordinate, $a$ the equilibrium lattice spacing, and $\{u_{n}\}$ the internal displacements. The lattice Hamiltonian separates exactly into CM and internal parts,
@@
where the hopping amplitudes depend linearly on the bond lengths,
@@
The electron-lattice coupling depends only on internal coordinates; the CM coordinate $R$ drops out exactly.
Strictly elastic backscattering via CM recoil
Because the full electron-lattice system is translationally invariant, there exists an exactly conserved total momentum. The electronic band energy satisfies $\varepsilon(k)=\varepsilon(-k)$, so reversing the electronic momentum does not change the electronic energy.
Consider an incoming electronic wavepacket narrowly peaked at momentum +k+k, and an internal lattice eigenstate $|\nu\rangle$. A scattering process of the form
$|k\rangle\otimes|\nu\rangle\otimes|P_{R}\rangle\;\longrightarrow\;|-k\rangle\otimes|\nu\rangle\otimes|P_{R}+2\hbar k\rangle$ (45)
leaves the internal phonon state unchanged and transfers the momentum $2\hbar k$ entirely to the CM of the chain.
@@
Exact energy conservation requires $\Delta E_{\rm CM}=0$, which is satisfied when
@@
For this value of the CM momentum, the backscattering process $k\rightarrow-k$ is strictly elastic, involving no excitation or absorption of internal phonon modes. Alternately, if $N$ is very large, the requirement $P_{R}=-\hbar k$ can be relaxed. For a finite, free chain of total mass $M$, electron backscattering $k\rightarrow-k$ changes the electron momentum by $\Delta p_{e}=-2\hbar k$. Momentum conservation then requires the chain to absorb a recoil $\Delta P_{\rm chain}=+2\hbar k$. If the chain initially has CM momentum $P$, energy conservation implies an associated recoil energy
$\Delta E_{\rm recoil}=\frac{(P+2\hbar k)^{2}-P^{2}}{2M}=\frac{2\hbar k\,(P+\hbar k)}{M}.$ (48)
Strictly elastic backscattering without internal excitation is therefore possible only for the special value $P=-\hbar k$; for generic initial CM momentum, an energy mismatch must be accommodated.
For a macroscopic chain, however, $M\propto N$ and the recoil energy scales as $\Delta E_{\rm recoil}\sim 1/N$, becoming parametrically small in the thermodynamic limit. In this regime the energy mismatch can be absorbed without resolving a discrete internal excitation. In particular, an electron wavepacket of spatial extent $L$ has an intrinsic energy spread
$\Delta E_{\rm pkt}\sim\hbar v_{g}\,\Delta k\sim\hbar v_{g}/L,$ (49)
which is independent of $N$ and typically exceeds $\Delta E_{\rm recoil}$ for sufficiently large chains. Equivalently, if the scattering occurs over a finite time $\tau_{\rm scatt}$, the associated time-energy uncertainty $\hbar/\tau_{\rm scatt}$ exceeds the recoil energy, rendering the process effectively elastic from the electron’s perspective.
In a phonon number-state description, the accommodation of this small energy mismatch corresponds to the creation or annihilation of one or more extremely low-frequency acoustic phonons. Importantly, however, such number-state changes do not account for the electron’s momentum reversal $k\rightarrow-k$: a single low-$q$ phonon carries momentum $\hbar q\ll 2\hbar k$. The dominant momentum transfer is instead absorbed by the collective motion of the chain (center-of-mass recoil), while the phonon sector merely provides an energetically soft channel required by exact energy conservation.
Thus, for large $N$, electron backscattering involves an exact momentum transfer of $2\hbar k$ to the background with a vanishing energy cost, and any phonon number-state changes reflect quasi-elastic energy bookkeeping rather than the mechanism of momentum reversal itself.
@@
with width $\sigma_{P}=\sqrt{M_{\rm tot}k_{B}T}$. The resonant value $P_{R}=-\hbar k$ therefore occurs with nonzero probability density for any $T>0$. Consequently, strictly elastic electronic backscattering events exist at finite temperature, mediated entirely by recoil of the lattice center of mass.
Bulk and boundary backscattering in a isolated chain
The elastic CM-recoil backscattering channel does not require the electron to encounter the ends of the chain and may occur locally in the bulk. The electron-lattice interaction in Eq. (44) depends only on nearest-neighbor bond distortions and is independent of absolute position.
For a wavepacket localized well away from the ends, with spatial extent $\ell\ll Na$, the matrix element connecting $|k\rangle$ and $|-k\rangle$ takes the form
$\langle-k,P_{R}+2\hbar k|H_{\rm int}|k,P_{R}\rangle\propto\sum_{n}e^{2ikna}\,\langle P_{R}+2\hbar k|P_{R}\rangle,$ (51)
where the sum runs over bulk sites only. The momentum transfer is absorbed entirely by the CM degree of freedom, while the internal lattice state remains unchanged.
Free boundaries of the finite chain can also reflect an electronic wavepacket once it reaches an end. Such boundary reflections are elastic and likewise involve a momentum reversal $k\rightarrow-k$ accompanied by CM recoil. Free boundaries therefore provide a spatial location for reflection but do not act as a momentum sink. Both bulk and boundary backscattering processes are governed by the same momentum conservation principle enforced by the CM zero mode.
@@
In addition to the strictly elastic CM recoil channel, a variety of inelastic and hybrid scattering processes are allowed within a isolated chain.
@@
If momentum is transferred to internal phonon modes, the internal lattice state changes. A generic single-phonon process takes the form
$|k\rangle\otimes|\{n_{q}\}\rangle\longrightarrow|k^{\prime}\rangle\otimes|\{n_{q}^{\prime}\}\rangle,$
with $n_{q}^{\prime}=n_{q}\pm 1$. Energy conservation requires
@@
The corresponding Golden-Rule rate is

### v2

@@ VII.4 Planckian diffusion and linear-in-$T$ resistivity
arising from elastic momentum scrambling in a time-dependent lattice background. In this formulation, diffusion emerges from quantum dynamics in a fluctuating potential and does not rely explicitly on strong inelastic scattering or rapid energy dissipation. Wave-on-wave simulations of the coupled electron–lattice problem demonstrate this mechanism explicitly [5, 16, 34].
@@
so that expressing $D$ in terms of an effective relaxation time yields [5]
@@
up to factors of order unity. From this viewpoint, the Planckian timescale may be understood not as a fundamental bound on inelastic scattering, but as an emergent parametrization of quantum-limited diffusion combined with equilibrium thermodynamics. It characterizes the rate at which charge spreads spatially rather than the rate at which energy must be irreversibly transferred to internal degrees of freedom.
@@ VIII Summary
In the conventional clamped-lattice formulation of electron–phonon scattering, the lattice is treated as a fixed reference frame and its translational zero mode does not appear explicitly among the dynamical variables. Within that representation, momentum transfer is naturally described in terms of internal phonon excitations. When the lattice center-of-mass degree of freedom is retained explicitly, an additional elastic momentum-transfer channel becomes manifest. In a finite crystal, pseudomomentum can be exchanged between an electron and the lattice background while total mechanical momentum is conserved through recoil of the lattice zero mode. Such processes do not require phonon creation and may contribute significantly to momentum relaxation at low and intermediate temperatures.
Making this channel explicit requires adopting a formulation in which (i) the lattice is treated as a finite, mechanically isolated object so that its center-of-mass motion remains dynamical, (ii) discrete translational symmetry is implemented symmetrically, with foreground and background pseudomomenta treated on equal footing, and (iii) the electron–lattice interaction is retained in its full density–density form. These steps do not alter the microscopic Hamiltonian; rather, they retain degrees of freedom that are commonly suppressed by choice of reference frame. When treated in this manner, elastic momentum exchange between electron and lattice background appears naturally within the same theoretical structure.
In this framework, pseudomomentum is redistributed between the electron and the lattice background, while total mechanical momentum is conserved through recoil of the translational zero mode. Momentum relaxation need not coincide with phonon creation, and energy dissipation need not set the transport rate. This separation of roles clarifies how efficient momentum randomization can coexist with comparatively slow energy relaxation.
Restoring explicit center-of-mass dynamics provides a more complete description of momentum bookkeeping in finite crystals. It offers a unified interpretation of experimental observations that indicate robust elastic scattering in clean materials, including weak localization, quantum oscillations, and ultrasonic attenuation. More broadly, the results suggest that in certain regimes, including some materials exhibiting linear-in-$T$ resistivity, momentum relaxation may be governed primarily by elastic scattering in a time-dependent lattice background [5], while inelastic processes play a secondary role in establishing global thermal equilibrium.
In a perfectly clean, isolated, translation-invariant crystal, total momentum is conserved exactly. Such a system may exhibit internal equilibration and decay of electronic current, but it cannot exhibit finite DC resistivity in the strict steady-state sense. True resistivity requires a mechanism for removing total momentum from the system.
IX Acknowledgments
@@
Figure 3: Isolated systems of mass $M+m$ containing a foreground particle of mass $m$ and a background of mass $M$. The particle is attracted to the background and moves frictionlessly, except in C and D it may excite vibrations of the chain. In A, the particle is free except for walls. In B, the particle is in a fixed periodic potential within its host, the box. In C we replace the fixed corrugation with a chain of masses connected by springs. In D, we dispense with the box, coming close to the situation that matters here. Of course the box could be sitting on a table, which could be fixed to a floating laboratory, which could be upon something so massive that it has significant gravitational attraction, yet that too is floating in space…
Appendix A Particle in a Free Box
We consider simple models that clarify the treatment of momentum. The usual particle-in-a-box problem fixes the walls and therefore violates momentum conservation. Here the box is given finite mass and allowed to move freely, so that total momentum is exactly conserved.
Consider a one-dimensional rigid box of mass $M$ and length $L$ with center-of-mass coordinate $R$, containing a particle of mass $m$ with coordinate $r$ (Fig. 3).
Introduce total and relative coordinates
$\mathcal{R}=\frac{mr+MR}{m+M},\qquad\rho=r-R,$
with total mass $M_{T}=m+M$ and reduced mass
$\mu=\frac{mM}{m+M}.$
The hard-wall constraint is $|\rho|<L/2$.
The Hamiltonian separates exactly:
$H=-\frac{\hbar^{2}}{2M_{T}}\frac{\partial^{2}}{\partial\mathcal{R}^{2}}+\left[-\frac{\hbar^{2}}{2\mu}\frac{\partial^{2}}{\partial\rho^{2}}+V_{\rm hw}(\rho)\right].$
Total momentum is conserved:
$\hat{P}_{\rm tot}=-i\hbar\frac{\partial}{\partial\mathcal{R}}.$
Case A: Flat interior
The relative Hamiltonian is
$H_{\rm rel}=-\frac{\hbar^{2}}{2\mu}\frac{d^{2}}{d\rho^{2}}+V_{\rm hw}(\rho).$
The normalized eigenfunctions are
$\phi_{n}(\rho)=\sqrt{\frac{2}{L}}\sin\!\left[\frac{n\pi}{L}\left(\rho+\frac{L}{2}\right)\right],\qquad n=1,2,\dots,$
with energies
$E_{n}=\frac{\hbar^{2}\pi^{2}n^{2}}{2\mu L^{2}}.$
Any internal state may be written
$\psi(\rho,t)=\sum_{n=1}^{\infty}c_{n}\,\phi_{n}(\rho)\,e^{-iE_{n}t/\hbar}.$
A Gaussian superposition centered at $\rho_{0}$ with mean wavenumber $k_{0}$,
$c_{n}=\mathcal{N}\exp\!\left[-\frac{(n-n_{0})^{2}}{4\sigma_{n}^{2}}\right]\exp\!\left[-ik_{n}\rho_{0}\right],\qquad k_{n}=\frac{n\pi}{L},$
produces a localized packet
$\psi(\rho,0)\approx A(\rho-\rho_{0})e^{ik_{0}(\rho-\rho_{0})}.$
Although each $\phi_{n}$ carries no relative momentum, the particle and box momenta may be nonzero individually, while $P_{\rm tot}$ remains fixed. At a wall collision both reverse sign, preserving total momentum.
Case B: Periodic interior
Let the interior potential be periodic relative to the box:
$V_{\rm per}(\rho)=V_{\rm per}(\rho+a).$
Then
$H_{\rm rel}=-\frac{\hbar^{2}}{2\mu}\frac{d^{2}}{d\rho^{2}}+V_{\rm per}(\rho)+V_{\rm hw}(\rho).$
Eigenfunctions take Bloch standing-wave form
$\varphi_{bn}(\rho)=\sqrt{\frac{2}{L}}u_{bn}(\rho)\sin\!\left(\frac{n\pi}{L}(\rho+L/2)\right),$
with
$u_{bn}(\rho+a)=u_{bn}(\rho).$
Wavepackets are superpositions
$\psi(\rho,0)=\sum_{n=1}^{\infty}A_{n}\varphi_{bn}(\rho).$
Shared pseudomomentum
Because $\rho=r-R$,
$e^{iq\rho}=e^{iqr}e^{-iqR}.$
Internal pseudomomentum $\hbar q$ is therefore shared between particle and box. For fixed total momentum $P_{0}$,
$p_{r}=\frac{m}{m+M}P_{0}+\hbar q,\qquad P_{R}=\frac{M}{m+M}P_{0}-\hbar q,$
so $p_{r}+P_{R}=P_{0}$.
Cases C and D: Dynamical periodic background
Let the periodic background be a dynamical elastic chain with atomic coordinates $R_{j}$:
$H=\frac{p_{r}^{2}}{2m}+\sum_{j}\frac{p_{j}^{2}}{2M}+\frac{K}{2}\sum_{j}(R_{j+1}-R_{j}-a)^{2}+H_{\rm int}.$
With density–density coupling
$H_{\rm int}=\int dr\,dr^{\prime}\,\rho_{p}(r)V(r-r^{\prime})\rho_{\rm lat}(r^{\prime}),$
where
$\rho_{p}(r)=\delta(r-\hat{r}),\qquad\rho_{\rm lat}(r)=\sum_{j}\delta(r-\hat{R}_{j}),$
the Hamiltonian depends only on coordinate differences and is invariant under rigid translation
$r\rightarrow r+a_{0},\qquad R_{j}\rightarrow R_{j}+a_{0}.$
Thus total momentum
$P_{\rm tot}=p_{r}+\sum_{j}p_{j}$
is exactly conserved. Diagonalizing the lattice,
$R_{j}=R_{\rm lat}+\sum_{q}u_{q}e^{iqja},$
reveals the zero mode $R_{\rm lat}$. The density operator contains the factor
$e^{-iqR_{\rm lat}},$
showing explicitly that momentum $\hbar q$ transferred to the particle is balanced by recoil of the lattice center of mass. Internal phonons redistribute momentum but are not required for momentum conservation.
Appendix B Elastic and inelastic scattering in a tight-binding, isolated one-dimensional chain
We consider a finite one-dimensional tight-binding chain of $N$ identical atoms, isolated in space, with nearest-neighbor springs and an electron moving in a basis of localized orbitals. The chain possesses an exact translational zero mode (center-of-mass motion). An electron can therefore exchange momentum with the lattice as a whole without necessarily exciting internal phonon modes. This opens a phonon-diagonal (internally elastic) backscattering channel for an electronic wavepacket, with the required momentum reversal balanced by center-of-mass recoil.
@@ Appendix B Elastic and inelastic scattering in a tight-binding, isolated one-dimensional chain
Write the atomic coordinates as
@@
where $R$ is the center-of-mass coordinate, $a$ the equilibrium lattice spacing, and $\{u_{n}\}$ the internal displacements. The lattice Hamiltonian separates into CM and internal parts,
@@
with bond-length-dependent hopping
@@
Thus the electron–lattice coupling depends only on internal coordinates $u_{n+1}-u_{n}$; the CM coordinate $R$ drops out exactly.
Phonon-diagonal backscattering via CM recoil
Because the full electron+lattice Hamiltonian is translationally invariant, the total momentum is exactly conserved. Since the band energy satisfies $\varepsilon(k)=\varepsilon(-k)$, electronic reversal $k\to-k$ does not change the electronic energy.
Consider an incoming wavepacket narrowly peaked at +k+k, an internal lattice eigenstate $|\nu\rangle$, and a CM momentum eigenstate $|P_{R}\rangle$. A phonon-diagonal backscattering channel has the form
$|k\rangle\otimes|\nu\rangle\otimes|P_{R}\rangle\longrightarrow|-k\rangle\otimes|\nu\rangle\otimes|P_{R}+2\hbar k\rangle,$ (45)
i.e. the internal state is unchanged while the CM absorbs the recoil $2\hbar k$.
@@
Exact energy conservation with strictly unchanged internal state requires $\Delta E_{\rm CM}=0$, i.e.
@@
For generic $P_{R}$ there is a recoil-energy mismatch
$\Delta E_{\rm recoil}=\frac{(P_{R}+2\hbar k)^{2}-P_{R}^{2}}{2M_{\rm tot}}=\frac{2\hbar k\,(P_{R}+\hbar k)}{M_{\rm tot}}.$ (48)
However, for a macroscopic chain $M_{\rm tot}\propto N$ so $|\Delta E_{\rm recoil}|\sim 1/N$ becomes parametrically small. For a wavepacket of spatial extent $L$ with intrinsic energy width $\Delta E_{\rm pkt}\sim\hbar v_{g}/L$, or for scattering occurring over a finite time $\tau_{\rm scatt}$ with uncertainty $\hbar/\tau_{\rm scatt}$, the process is effectively elastic whenever $\Delta E_{\rm recoil}$ is below the available energy resolution. In a number-state phonon description, any residual mismatch may be accommodated by very soft internal modes, but the momentum reversal $k\to-k$ is carried by CM recoil; internal excitations, if present, serve primarily as an energetically low-cost bookkeeping channel.
@@
with width $\sigma_{P}=\sqrt{M_{\rm tot}k_{B}T}$. The resonant value $P_{R}=-\hbar k$ therefore occurs with nonzero probability density for any $T>0$, so strictly phonon-diagonal backscattering events exist at finite temperature in an isolated chain.
Bulk and boundary backscattering
The CM-recoil channel does not require the electron to reach the ends: it can occur in the bulk from the static component of thermal distortions (phonon-diagonal scattering). The SSH coupling
$H_{\rm int}=\alpha\sum_{n=1}^{N-1}(u_{n+1}-u_{n})\left(c_{n+1}^{\dagger}c_{n}+c_{n}^{\dagger}c_{n+1}\right)$
connects $|k\rangle$ and $|-k\rangle$ through the $2k$ Fourier component of the bond-distortion field. CM recoil is enforced by the CM translation factor. In particular, the CM shift operator produces
$\langle P_{R}^{\prime}|e^{-i(2k)R}|P_{R}\rangle\propto\delta(P_{R}^{\prime}-P_{R}-2\hbar k),$ (50)
so momentum conservation is implemented by a delta function (or Kronecker delta in a finite quantization volume), not by an inner product between distinct CM momentum eigenstates.
Free boundaries can also reflect a wavepacket once it reaches an end. Such reflections reverse $k$ and are likewise accompanied by CM recoil. Boundaries provide a location for reflection but do not constitute a momentum sink; both bulk (distortion-mediated) and boundary reflections obey the same conservation principle enforced by the CM zero mode.
@@
In addition to the phonon-diagonal CM-recoil channel, inelastic and hybrid processes are allowed.
@@
If momentum is transferred to internal modes, the internal lattice state changes:
$|k\rangle\otimes|\{n_{q}\}\rangle\to|k^{\prime}\rangle\otimes|\{n_{q}^{\prime}\}\rangle,\qquad n_{q}^{\prime}=n_{q}\pm 1.$
Energy conservation requires
@@
A Golden-Rule rate has the form